OpenStudy (anonymous):

If x is -4 and y is 5 then how do I solve x2+2xy+y2

OpenStudy (anonymous):

What you want to do is substitute -4 for x and 5 for y (-4)^2 + 2(-4)(5) + (5)^2 Are you with me so far?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Then you can simplify 16 + -20 + 25 = 16 + 5 =21

OpenStudy (anonymous):

Ok I see now Ty

OpenStudy (loser66):

@KirbyLegs Question: What's wrong with my logic?? Please, state out x^2 + 2xy +y^2 = (x+y)^2 replace x = -4, y =5 (-4+5)^2 = 1